Write the system of linear equations represented by the following matrix. Use x, y, & z as variables.
[7 0 4 |-13]
[0 1 -5 | 11]
[2 7 0 | 6]
Thanks in advance

Answer by jim_thompson5910(35256)

[7 0 4 |-13]
[0 1 -5 | 11]
[2 7 0 | 6]

converts to

7x+0y+4z = -13
0x+1y-5z = 11
2x+7y+0z = 6

which simplifies to

7x+4z = -13
y-5z = 11
2x+7y = 6